OpenStudy (anonymous):

Solve for x. y=(2-6k)x I need to know how to do this step by step please HELP!!!!

jimthompson5910 (jim_thompson5910):

Since you're multiplying x by 2-6k, you can "undo" this by dividing both sides by 2-6k to get \[\Large \frac{y}{2-6k}=x\] then flipping both sides to get \[\Large x=\frac{y}{2-6k}\]

OpenStudy (anonymous):

Thanks hun so much!!!! Wow that was so easy lol!

OpenStudy (anonymous):

Now I understand once I saw how you did it